OpenStudy (calculusxy):

MEDAL!!! Find the sum of the arithmetic series described. d = 2, n = 26, \(a_n\) = 42

OpenStudy (calculusxy):

@welshfella @zepdrix

OpenStudy (welshfella):

hmmm i think you need to find the first term a1 = 42 - 25(2) = -10 so S26 = (26/2) [-10 + 42]

OpenStudy (calculusxy):

Why did you make that equation for finding the first term?

OpenStudy (welshfella):

well i intended to use Sn = (n/2) [ a1 + l] where l = last term

OpenStudy (welshfella):

not sure if that was is right .....

OpenStudy (calculusxy):

can i use the explicit formula for finding the first term?

OpenStudy (welshfella):

No - its -8 I cant count!!

OpenStudy (calculusxy):

2(25) + a1 = 42 a1= -8

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

now you can use (26/2) [ -8 + 42]

OpenStudy (calculusxy):

i got 442

OpenStudy (welshfella):

yea

OpenStudy (calculusxy):

thank you! can you help me with two more questions?

OpenStudy (welshfella):

i cant I have to go right now

OpenStudy (calculusxy):

ok thanks for your help though!